# Entire Large Solutions for Competitive Semilinear Elliptic Systems with General Nonlinearities Satisfying Keller–Osserman Conditions

### v2

@@ Abstract
We generalize a theorem of Lair concerning the existence of positive entire large solutions to competitive semilinear elliptic systems. While Lair’s original result [3] was established for power-type nonlinearities, our work extends the theory to a broad class of general nonlinearities satisfying a Keller–Osserman-type growth condition. The proof follows the same conceptual framework monotone iteration to construct global positive solutions, reduction to a scalar inequality for the sum of the components, application of a Keller–Osserman transform, and a two-step radial integration argument but replaces the explicit power-law growth with a general monotone envelope function. This approach yields a unified and verifiable criterion for the existence of large solutions in terms of the Keller–Osserman integral, thereby encompassing both critical and supercritical growth regimes within a single analytical setting.
@@ 1 Introduction
The search for positive entire large solutions to nonlinear elliptic systems is a classical theme in analysis, with deep connections to geometry, physics, and population models. In this work, we focus on the radial semilinear system
@@
Our aim is to generalize the existence theory of Lair [3] "originally for power-type nonlinearities" to a broad class of nonlinearities governed by the Keller–Osserman growth condition. This unifies the treatment of critical and supercritical regimes, and recovers as special cases the results of Keller [2], Osserman [5], Cîrstea and Rădulescu [1], among others.
We call $(u,v)$ a positive entire solution if it is $C^{2}$, radial, and strictly positive on $0,\infty)$. The set of central values
@@
encodes the initial data leading to such solutions. We show that $\mathcal{G}$ has a monotone structure, certain subsets are bounded and closed, and that on its boundary, all solutions are large at least one component diverges as $r\rightarrow\infty$.
Our method adapts Lair’s barrier and radial integration techniques [4, 3], replacing explicit power laws with a general monotone envelope $h$ satisfying the Keller–Osserman condition. This approach not only recovers known power-type results but also links naturally to complementary bounded-solution theories [8, 6, 7].
Structure of the paper. Section 2 states the main results. Section 3 contains the proofs, built on monotone iteration, Keller–Osserman estimates, and barrier arguments. Section 4 concludes with remarks and perspectives.
@@ 2 Main Results
[…]
Figure 1: Existence region in the $(\alpha,\beta)$-plane: $G$ (green), $G_{\delta}$ (yellow), and $E(G_{\delta})$ (blue).
@@ 3 Proof of the Theorems
The proof of Theorem 2.1 is organized into six logically steps. We adapt the monotone iteration method of [3] to the present setting, where $f$ and $g$ satisfy the general hypotheses (H2)–(H3).
Step 1: Iterative scheme.
@@
Given $(\alpha_{0},\beta_{0})\in(0,\infty)\times(0,\infty)$, define
@@
Step 2: Monotonicity.
Proof.
By the positivity of $p,q$ and the monotonicity of $f,g$ in each variable (assumption (H2)), we have
@@
Thus, for each fixed $r$, both sequences $\{u_{k}(r)\}$ and $\{v_{k}(r)\}$ are nondecreasing. ∎
Step 3: Uniform upper bound via a scalar supersolution.
Proof.
Construction of $w$. The Keller–Osserman condition (H3) ensures the existence of a scalar function $w\in C^{2}([0,\infty))$ satisfying
@@
To construct $w$, consider the scalar radial ODE
$W^{\prime\prime}(r)+\frac{n-1}{r}W^{\prime}(r)=(p(r)+q(r))\,S(W(r)),$
where $S(s):=f(s,s)+g(s,s)$. By (H2)–(H3), $H$ is continuous, nondecreasing, positive for $s>0$, and satisfies the Keller–Osserman integral condition
$\int_{1}^{\infty}\frac{ds}{\sqrt{\int_{0}^{s}S(\tau)\,d\tau}}<\infty.$
Choosing $W(0)$ sufficiently large, standard ODE theory and the above integral condition guarantee that $W(r)\rightarrow\infty$ as $r\rightarrow\infty$. Setting $w:=W$ yields (3.3).
@@
  • It ensures that the iteration cannot “blow up” prematurely on finite intervals.
By the comparison principle for radial integral equations, we have
@@
Step 4: Convergence of the iteration.
@@
For each fixed $r\geq 0$, the sequences $\{u_{k}(r)\}$ and $\{v_{k}(r)\}$ are monotone increasing and bounded above by $w(r)$. Thus, the pointwise limits
exist for all $r\geq 0$.
Monotone Convergence Theorem (MCT). The MCT states: If $\{s_{k}\}$ is a sequence of nonnegative measurable functions such that $s_{k}(x)\uparrow s(x)$ pointwise, then
$\lim_{k\rightarrow\infty}\int s_{k}=\int\lim_{k\rightarrow\infty}s_{k}.$
In our setting, $s_{k}$ corresponds to the integrands in (3.1)–(3.2). Since $u_{k},v_{k}$ increase pointwise and $p,q,f,g$ are nonnegative, the integrands are nonnegative and increasing in $k$. Thus, MCT justifies passing the limit $k\rightarrow\infty$ inside the integrals, showing that $(u,v)$ satisfies (3.1)–(3.2). ∎
Step 5: Conclusion.
Proof.
The limit pair $(u,v)$ is a positive $C^{2}$ radial solution of (1.1) with $(u(0),v(0))=(\alpha_{0},\beta_{0})$. Thus $(\alpha_{0},\beta_{0})\in\mathcal{G}$, proving that $\mathcal{G}$ is nonempty. ∎
Step 6: Downward closure property.
Proof.
If $(\alpha_{0},\beta_{0})\in\mathcal{G}$ and $0<\alpha<\alpha_{0}$, $0<\beta<\beta_{0}$, then starting the iteration from $(\alpha,\beta)$ yields smaller initial data. By the same monotonicity and boundedness arguments, the iteration converges to a positive entire solution, hence $(\alpha,\beta)\in\mathcal{G}$. ∎
3.2 Proof of Theorem 2.2
@@
Part I: Boundedness of $\mathcal{G}_{\delta}$
Proof.
@@
Therefore, $\mathcal{G}_{\delta}$ cannot be unbounded; it is bounded. ∎
Part II: Closedness of $\mathcal{G}_{\delta}$
Proof.
@@
Step 2: Uniform bounds and equicontinuity. From the monotone iteration scheme in Theorem 2.1 and the Keller–Osserman supersolution $w$ from (3.3) with $w(0)>\alpha_{m}+\beta_{m}$, we deduce that the family $\{(u_{m},v_{m})\}$ is uniformly bounded on compact intervals and equicontinuous.
Step 3: Compactness and passage to the limit. By the Arzelá –Ascoli theorem, there exists a subsequence converging uniformly on compact sets to a limit $(u,v)$. Passing to the limit in (3.1)–(3.2) (justified by the dominated convergence theorem, since the integrands are continuous and uniformly bounded on compacts) shows that $(u,v)$ is a positive entire solution with
$u(0)=\alpha,\quad v(0)=\beta.$
Since $\min\{\alpha,\beta\}\geq\delta$ by construction, we have $(\alpha,\beta)\in\mathcal{G}_{\delta}$.
Conclusion: $\mathcal{G}_{\delta}$ is both bounded and closed. ∎
@@
For clarity, we structure the proof of Theorem 2.3 into seven distinct and logically self-contained steps.
Standing assumptions and notation.
@@
We work under hypotheses (H1)–(H3) and consider the radial system (1.1) in its equivalent integral (Green) form (3.1)–(3.2), with prescribed central values
@@
  • $H^{\prime\prime}(s)\geq 0$ since $s\mapsto f(s,s)+g(s,s)$ is nondecreasing by (H2).
Step 1: Perturbation of initial data and blow-up for approximants
Proof.
From Theorem 2.1, $\mathcal{G}$ is nonempty and downward closed in each coordinate. From Theorem 2.2, $\mathcal{G}_{\delta}$ is closed and bounded for $\delta>0$ sufficiently small.
@@
Step 2: Reduction to a scalar inequality
Proof.
Let
@@
by (H2). ∎
Step 3: Applying the Keller–Osserman transform
Proof.
Applying $H$ to $W_{k}$ and using $H^{\prime\prime}\geq 0$, the chain rule gives
$\Delta H(W_{k})=H^{\prime}(W_{k})\,\Delta W_{k}+H^{\prime\prime}(W_{k})|\nabla W_{k}|^{2}\geq H^{\prime}(W_{k})\,\Delta W_{k}.$
@@
Step 4: First radial integration
Proof.
In radial coordinates, the above inequality becomes
@@
$(H(W_{k}))^{\prime}\geq-r^{1-n}\int_{0}^{r}s^{n-1}(p(s)+q(s))\,ds.$
Step 5: Second radial integration and lower bound
Proof.
Integrating from $r$ to $R_{k}$ gives
@@
If $R<\infty$, the right-hand side tends to $\infty$ as $r\rightarrow R^{-}$, contradicting the fact that $(u,v)$ is entire. Hence $R=\infty$, and we have established the quantitative lower bound (2.2). ∎
Step 6: Divergence of at least one component.
Proof.
From the previous steps (in particular, Step 1 of the proof of Theorem 2.2 and (3.4)), we already know that
@@
Proof of Claim 2. The argument is symmetric, using $\ell_{g}$, $q$, and a barrier $w$ with $w(R_{1})<v(R_{1})$ to obtain the same type of contradiction. ∎∎
Step 7: Conclusion.
Proof.
We have now completed the construction of a positive, radial, entire solution
@@
$W(r)\longrightarrow+\infty\quad asr\rightarrow\infty,$

### v3

@@ Abstract
We establish the existence of positive entire large solutions for competitive semilinear elliptic systems with general nonlinearities, significantly extending recent work of Lair [3] from power-type to Keller–Osserman-type growth conditions. For the radial system $\Delta u=p(r)f(u,v)$, $\Delta v=q(r)g(u,v)$ in $\mathbb{R}^{n}$ ($n\geq 3$), we prove that under appropriate integrability conditions on the weights $p,q$ and a Keller–Osserman condition on the envelope function $h$ satisfying $f(s,t)+g(s,t)\geq h(s+t)$, there exists a positive radial solution $(u,v)$ with at least one component diverging to infinity. Our proof employs monotone iteration, scalar reduction via comparison principles, the Keller–Osserman transform, and barrier arguments. The main result provides explicit quantitative lower bounds on the blow-up rate and establishes conditions under which both components diverge simultaneously. This unified framework recovers classical power-law results as special cases while encompassing critical and supercritical growth regimes, and complements the bounded-solution theory in the literature.
@@ 1 Introduction
The study of positive entire large solutions to nonlinear elliptic systems is a fundamental area in the theory of partial differential equations, with profound connections to differential geometry, mathematical physics, reaction-diffusion systems, and population dynamics. Solutions that blow up at infinity arise naturally in various contexts: prescribed scalar curvature problems [2, 5], conformal geometry, steady-state models of chemical reactions, and competitive or cooperative biological systems.
The existence and qualitative behavior of such large solutions have been investigated extensively since the pioneering works of Keller [2] and Osserman [5] in the 1950s, who established the now-classical Keller–Osserman condition for scalar equations. This condition characterizes when a single semilinear equation $\Delta u=f(u)$ admits solutions that diverge to infinity at spatial infinity. Extensions to systems began with the work of Cîrstea and Rădulescu [1], and have seen significant developments in recent years, including the work of Zhang et al. [8] on bounded solutions, Peterson and Wood [6] on non-monotone systems, and more recently, Lair [3] for competitive systems with power-type nonlinearities.
In this work, we focus on the radial semilinear system
@@
in the entire space $\mathbb{R}^{n}$, where the weight functions $p,q$ are radially symmetric and satisfy appropriate integrability conditions, and the nonlinearities $f,g$ satisfy general Keller–Osserman-type growth conditions.
@@
Main contributions. Our principal aim is to extend the existence theory of Lair [3], originally established for power-type nonlinearities, to a broad class of nonlinearities satisfying Keller–Osserman-type growth conditions. This extension is both natural and necessary:
  • (i)
    Generality: Many physical and geometric problems lead to nonlinearities that do not have pure power-law form. Our framework accommodates these more general growth conditions while preserving the essential analytical structure.
  • (ii)
    Unification: By working with a general envelope function $h$ satisfying the Keller–Osserman integral condition, we provide a unified treatment of critical and supercritical growth regimes, recovering as special cases not only Lair’s results but also classical results of Keller [2], Osserman [5], and Cîrstea and Rădulescu [1].
  • (iii)
    Completeness: Our results complement the bounded-solution theories developed by Zhang et al. [8], Peterson and Wood [6], and Santos et al. [7], providing a complete picture of the existence landscape for system (1.1).
Analytical framework. We call $(u,v)$ a positive entire solution if it is $C^{2}$, radial, and strictly positive on $0,\infty)$. A solution is called large if at least one component diverges to infinity as $r\to\infty$. The set of central values
@@
encodes the initial data leading to such solutions. A key insight of our work is that $\mathcal{G}$ has a rich geometric structure: it is downward-closed (in the sense that smaller initial values also lead to entire solutions), certain regularized subsets $\mathcal{G}_{\delta}$ are compact, and solutions corresponding to boundary points of $\mathcal{G}_{\delta}$ are necessarily large—at least one component diverges as $r\rightarrow\infty$.
Methodology. Our approach adapts and extends Lair’s barrier and radial integration techniques [4, 3] to the setting of general nonlinearities. The key steps are:
  • Monotone iteration: We construct solutions via a monotone iterative scheme, with upper bounds provided by scalar supersolutions.
  • Scalar reduction: By exploiting the monotonicity of $f$ and $g$, we reduce the coupled system to a scalar differential inequality for $W=u+v$.
  • Keller–Osserman transform: We apply a nonlinear transformation $H$ to $W$, converting the nonlinear problem into a linear inequality amenable to explicit integration.
  • Blow-up dichotomy: Using barrier arguments and the Hopf lemma, we establish that solutions with boundary initial data must blow up.
This approach not only recovers known power-type results but also provides explicit quantitative bounds on the blow-up rate.
Organization. Section 2 states the main results. Section 3 contains the complete proofs, built on monotone iteration, Keller–Osserman estimates, and barrier arguments. Section 4 concludes with remarks and perspectives for future research.
@@ 2 Main Results
$\alpha$$\beta$$\alpha=\delta$$\beta=\delta$$\mathcal{G}$$\mathcal{G}_{\delta}$$\mathcal{E}(\mathcal{G}_{\delta})$
Figure 1: Existence region in the $(\alpha,\beta)$-plane: $\mathcal{G}$ (green), $\mathcal{G}_{\delta}$ (yellow), and $\mathcal{E}(\mathcal{G}_{\delta})$ (blue).
@@ 3 Proof of the Theorems
The proof is organized into six logical steps. We adapt the monotone iteration method of [3] to the present setting, where $f$ and $g$ satisfy the general hypotheses (H2)–(H3).
Step 1: Iterative scheme. Given $(\alpha_{0},\beta_{0})\in(0,\infty)\times(0,\infty)$, define
@@
Step 2: Monotonicity. By the positivity of $p,q$ and the monotonicity of $f,g$ in each variable (assumption (H2)), we have
@@
Thus, for each fixed $r$, both sequences $\{u_{k}(r)\}$ and $\{v_{k}(r)\}$ are nondecreasing.
Step 3: Uniform upper bound via a scalar supersolution. Construction of $w$. The Keller–Osserman condition (H3) ensures the existence of a scalar function $w\in C^{2}([0,\infty))$ satisfying
@@
To construct $w$ rigorously, consider the scalar radial ODE
$W^{\prime\prime}(r)+\frac{n-1}{r}W^{\prime}(r)=(p(r)+q(r))\,S(W(r)),$ (3.4)
where $S(s):=f(s,s)+g(s,s)$. By (H2)–(H3), $S$ is continuous, nondecreasing, positive for $s>0$, vanishes at $s=0$, and satisfies the Keller–Osserman integral condition
$\int_{1}^{\infty}\frac{ds}{\sqrt{\int_{0}^{s}S(\tau)\,d\tau}}<\infty.$ (3.5)
By classical existence theory for radial ODEs (see [2, 5]), for any initial value $W(0)=W_{0}>0$, there exists a maximal solution $W$ of (3.4) on an interval $0,R_{W_{0}})$ with $R_{W_{0}}\in(0,\infty]$. The Keller–Osserman condition (3.5) guarantees that if $W_{0}$ is chosen sufficiently large, then $R_{W_{0}}=\infty$ and $W(r)\to\infty$ as $r\to\infty$ (see [1], Theorem 2.1). Specifically, choosing $W_{0}>\alpha_{0}+\beta_{0}$, we obtain a global solution $W$ satisfying $W(0)=W_{0}>\alpha_{0}+\beta_{0}$ and $\lim_{r\to\infty}W(r)=\infty$. Setting $w:=W$ yields (3.3).
@@
  • It ensures that the iteration cannot blow up prematurely on finite intervals.
By the comparison principle for radial integral equations (which follows from the maximum principle for the Laplacian), since $u_{0}(r)+v_{0}(r)=\alpha_{0}+\beta_{0}<w(0)$ and
$\Delta(u_{k}+v_{k})=p(r)f(u_{k},v_{k})+q(r)g(u_{k},v_{k})\leq(p(r)+q(r))\big(f(u_{k}+v_{k},u_{k}+v_{k})+g(u_{k}+v_{k},u_{k}+v_{k})\big),$
by monotonicity of $f,g$, we deduce by induction that
@@
Step 4: Convergence of the iteration. For each fixed $r\geq 0$, the sequences $\{u_{k}(r)\}$ and $\{v_{k}(r)\}$ are monotone increasing (by Step 2) and bounded above by $w(r)$ (by Step 3). Thus, the pointwise limits
exist for all $r\geq 0$ by monotone convergence of real sequences.
Justification via the Monotone Convergence Theorem (MCT). To verify that $(u,v)$ satisfies the integral system (3.1)–(3.2), we apply Lebesgue’s MCT. The MCT states: If $\{\phi_{k}\}$ is a sequence of nonnegative measurable functions such that $\phi_{k}(x)\uparrow\phi(x)$ pointwise almost everywhere, then
$\lim_{k\rightarrow\infty}\int\phi_{k}\,d\mu=\int\lim_{k\rightarrow\infty}\phi_{k}\,d\mu=\int\phi\,d\mu.$
In our setting, fix $r>0$ and consider the innermost integral in (3.1):
$I_{k}(t):=\int_{0}^{t}s^{n-1}p(s)f(u_{k-1}(s),v_{k-1}(s))\,ds.$
Since $u_{k-1}(s)\uparrow u(s)$ and $v_{k-1}(s)\uparrow v(s)$ pointwise, and $f$ is nondecreasing and continuous, we have
$p(s)f(u_{k-1}(s),v_{k-1}(s))\uparrow p(s)f(u(s),v(s))\quad\text{pointwise in }s.$
The integrands are nonnegative. By MCT,
$\lim_{k\to\infty}I_{k}(t)=\int_{0}^{t}s^{n-1}p(s)f(u(s),v(s))\,ds.$
Moreover, $I_{k}(t)\uparrow$ and is bounded by
$I_{k}(t)\leq\int_{0}^{t}s^{n-1}(p(s)+q(s))S(w(s))\,ds=:M(t)<\infty$
for each $t<\infty$, by construction of $w$. Applying MCT again to the outer integral in (3.1), we obtain
$\lim_{k\to\infty}u_{k}(r)=\alpha_{0}+\int_{0}^{r}t^{1-n}\left(\lim_{k\to\infty}I_{k}(t)\right)dt=\alpha_{0}+\int_{0}^{r}t^{1-n}\int_{0}^{t}s^{n-1}p(s)f(u(s),v(s))\,ds\,dt.$
Thus $u=\lim_{k\to\infty}u_{k}$ satisfies (3.1). The same argument applies to $v$. By standard elliptic regularity, $(u,v)\in C^{2}([0,\infty))\times C^{2}([0,\infty))$.
Step 5: Conclusion. The limit pair $(u,v)$ is a positive $C^{2}$ radial solution of (1.1) with $(u(0),v(0))=(\alpha_{0},\beta_{0})$. Thus $(\alpha_{0},\beta_{0})\in\mathcal{G}$, proving that $\mathcal{G}$ is nonempty.
Step 6: Downward closure property. If $(\alpha_{0},\beta_{0})\in\mathcal{G}$ and $0<\alpha<\alpha_{0}$, $0<\beta<\beta_{0}$, then starting the iteration from $(\alpha,\beta)$ yields smaller initial data. By the same monotonicity and boundedness arguments (Steps 1–4), the iteration converges to a positive entire solution with initial values $(\alpha,\beta)$, hence $(\alpha,\beta)\in\mathcal{G}$. This completes the proof of Theorem 2.1. ∎
3.2 Proof of Theorem 2.2
@@
Part I: Boundedness of $\mathcal{G}_{\delta}$.
@@
Therefore, $\mathcal{G}_{\delta}$ cannot be unbounded; it is bounded.
Part II: Closedness of $\mathcal{G}_{\delta}$.
@@
Step 2: Uniform bounds and equicontinuity. From the monotone iteration scheme in Theorem 2.1 and the Keller–Osserman supersolution $w$ from (3.3) with $w(0)>\sup_{m}(\alpha_{m}+\beta_{m})$, we deduce that
$u_{m}(r)+v_{m}(r)\leq w(r)\quad\forall r\geq 0,\,\forall m\geq 1.$
Hence the family $\{u_{m}\}$ and $\{v_{m}\}$ is uniformly bounded on each compact interval $[0,R]$.
To establish equicontinuity, we use the integral representation (3.1)–(3.2). For $r_{1},r_{2}\in[0,R]$ with $r_{1}<r_{2}$, we have
$|u_{m}(r_{2})-u_{m}(r_{1})|$ $\leq\int_{r_{1}}^{r_{2}}t^{1-n}\int_{0}^{t}s^{n-1}p(s)f(u_{m}(s),v_{m}(s))\,ds\,dt$
$\leq\int_{r_{1}}^{r_{2}}t^{1-n}\int_{0}^{t}s^{n-1}(p(s)+q(s))S(w(s))\,ds\,dt,$
where $S(s)=f(s,s)+g(s,s)$. Since the right-hand side is independent of $m$ and tends to zero as $r_{2}-r_{1}\to 0$ (by absolute continuity of the integral), the family $\{u_{m}\}$ is equicontinuous on $[0,R]$. The same argument applies to $\{v_{m}\}$.
Step 3: Compactness and passage to the limit. By the Arzelà–Ascoli theorem, the sequence $\{(u_{m},v_{m})\}$ is relatively compact in $C([0,R])\times C([0,R])$ for each $R>0$. By a standard diagonal argument, there exists a subsequence $(u_{m_{j}},v_{m_{j}})$ converging uniformly on every compact interval $[0,R]$ to a limit $(u,v)\in C([0,\infty))\times C([0,\infty))$.
To verify that $(u,v)$ satisfies (3.1)–(3.2), fix $r>0$ and observe that the integrands in (3.1) converge:
$s^{n-1}p(s)f(u_{m_{j}}(s),v_{m_{j}}(s))\to s^{n-1}p(s)f(u(s),v(s))$
pointwise in $s\in[0,r]$ by continuity of $f$, and are uniformly bounded by $s^{n-1}(p(s)+q(s))S(w(s))\in L^{1}([0,r])$. By the dominated convergence theorem,
$\lim_{j\to\infty}\int_{0}^{t}s^{n-1}p(s)f(u_{m_{j}}(s),v_{m_{j}}(s))\,ds=\int_{0}^{t}s^{n-1}p(s)f(u(s),v(s))\,ds$
uniformly in $t\in[0,r]$. Applying dominated convergence again to the outer integral in (3.1), we conclude that $(u,v)$ satisfies (3.1)–(3.2), hence is a positive entire solution with
$u(0)=\lim_{j\to\infty}u_{m_{j}}(0)=\lim_{j\to\infty}\alpha_{m_{j}}=\alpha,\quad v(0)=\beta.$
Since $\min\{\alpha,\beta\}\geq\delta$ by construction (as the limit of a sequence in $\mathcal{G}_{\delta}$), we have $(\alpha,\beta)\in\mathcal{G}_{\delta}$.
Conclusion: $\mathcal{G}_{\delta}$ is both bounded and closed, completing the proof of Theorem 2.2. ∎
@@
We structure the proof into seven distinct and logically self-contained steps.
Standing assumptions and notation. We work under hypotheses (H1)–(H3) and consider the radial system (1.1) in its equivalent integral (Green) form (3.1)–(3.2), with prescribed central values
@@
  • $H^{\prime\prime}(s)\geq 0$ since $s\mapsto f(s,s)+g(s,s)$ is nondecreasing by (H2). Explicitly,
    $H^{\prime\prime}(s)=\frac{d}{ds}\left(-\frac{1}{f(s,s)+g(s,s)}\right)=\frac{\frac{\partial}{\partial s}[f(s,s)+g(s,s)]}{[f(s,s)+g(s,s)]^{2}}\geq 0.$
Step 1: Perturbation of initial data and blow-up for approximants. From Theorem 2.1, $\mathcal{G}$ is nonempty and downward closed in each coordinate. From Theorem 2.2, $\mathcal{G}_{\delta}$ is closed and bounded for $\delta>0$ sufficiently small.
@@
Step 2: Reduction to a scalar inequality. Let
@@
by (H2).
Step 3: Applying the Keller–Osserman transform. Applying $H$ to $W_{k}$ and using $H^{\prime\prime}\geq 0$, the chain rule gives
$\Delta H(W_{k})=H^{\prime}(W_{k})\,\Delta W_{k}+H^{\prime\prime}(W_{k})|\nabla W_{k}|^{2}\geq H^{\prime}(W_{k})\,\Delta W_{k}.$ (3.8)
Justification of the chain rule: Since $W_{k}\in C^{2}$ and $H\in C^{2}((0,\infty))$, the composition $H\circ W_{k}$ is $C^{2}$ wherever $W_{k}>0$. The Laplacian of the composition is
$\Delta H(W_{k})=H^{\prime}(W_{k})\Delta W_{k}+H^{\prime\prime}(W_{k})|\nabla W_{k}|^{2},$
which is the standard chain rule for the Laplacian. Since $H^{\prime\prime}(W_{k})\geq 0$ by our earlier computation, we obtain (3.8).
@@
Step 4: First radial integration. In radial coordinates, the above inequality becomes
@@
$(H(W_{k}))^{\prime}(r)\geq-r^{1-n}\int_{0}^{r}s^{n-1}(p(s)+q(s))\,ds.$
Step 5: Second radial integration and lower bound. Integrating from $r$ to $R_{k}$ gives
@@
If $R<\infty$, the right-hand side tends to $\infty$ as $r\rightarrow R^{-}$, contradicting the fact that $(u,v)$ is entire. Hence $R=\infty$, and we have established the quantitative lower bound (2.2).
Step 6: Divergence of at least one component. From the previous steps (in particular, Step 1 of the proof of Theorem 2.2 and (3.6)), we already know that
@@
Proof of Claim 2. The argument is symmetric, using $\ell_{g}$, $q$, and a barrier $w$ with $w(R_{1})<v(R_{1})$ to obtain the same type of contradiction. ∎
Step 7: Conclusion. We have now completed the construction of a positive, radial, entire solution
@@
$W(r)\longrightarrow+\infty\quad\text{as }r\rightarrow\infty,$